# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v2

@@ III.5 Polycatenanes: Extensional Flow
Fig. 6a shows that the variation of fractional extension of polycatenanes in the flow direction is less than linear polymers, and it is lower in polycatenanes that have fewer rings. For instance, a polycatenane with 8 beads per ring and 5 rings extends only 2.4 times beyond equilibrium, while a polycatenane chain with the same ring size and 65 rings stretches 4.2 times. More rings give the chain more degrees of freedom, leading to greater relative fluctuations in size and shape. In the limiting case, if we continue increasing the number of rings indefinitely, the chain approaches a linear polymer Brereton (2001). At high Wi, the difference becomes smaller as both polycatenanes and linear polymers approach fully extended configurations. As shown in Fig. 6b, increasing the ring size while keeping the number of rings constant also leads to more variability in extension due to the higher conformational degrees of freedom from the larger rings. The chains with larger rings have more flexibility Rauscher et al. (2020) and can collapse and extend more readily. As a result, in weak flows, chains with smaller rings exhibit higher $\langle X_{bb}\rangle/L$, whereas in strong flows, chains with larger rings show greater extension.
@@
Extensional viscosity ($\eta_{e}$) is significantly lower in polycatenanes compared to linear polymers with the same molecular weight, as shown in Fig. 6e. For instance, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) has 22 times lower ($\eta_{e}$) than a linear polymer with 80 beads at $\mathrm{Wi}=10$. Furthermore, adding more rings to a polycatenane increases the extensional viscosity due to increased molecular weight. The same effect is achieved by making the rings larger, as seen in Fig. 6f. Comparing a polycatenane with 16 beads per ring and 10 rings ($M_{w}=160$) to a polycatenane with 8 beads per ring and 21 rings ($M_{w}=168$) allows us to decouple the effects of molecular weight from the impact of chain topology on viscosity. Based on Figs. 6e and 6f, at low Wi, the chain with more small rings has higher viscosity. This reflects the more extended conformation of the smaller rings, leading to higher stress and resistance to extension. At high Wi, the chain with fewer large rings show significantly higher extensional viscosity, up to 270% higher at the highest Wi. Under strong extensional flow, the rings in the chain with fewer large rings are more effectively stretched, leading to an increased resistance to further deformation. The extended conformation under high strain produces a larger stress response, which is reflected in the higher measured extensional viscosity. In contrast, the greater number of flexible mechanical interlocks in the chain with more small rings facilitates lower stress response when subjected to high strain rates which ultimately reduces $\eta_{e}$ relative to the chain with fewer large rings.
@@ III.6 Polycatenanes: Shear Flow
As shown in Fig. 7a, at low Wi, polycatenanes exhibit a range of extensional behaviors based on the number of rings in their structure. Chains with fewer rings have limited variation of $\langle X_{bb}\rangle/L$ throughout the Wi range, while chains with higher number of rings can be deformed more readily in the flow, leading to larger variation of fractional extension in flow direction. For example, a polycatenane of 5 rings stretches between 0.32 and 0.43 of its maximum, compared to a 65 ring polycatenane, which extends between 0.17 and 0.35. Shorter, less flexible polycatenanes begin to stretch in the flow direction at lower flow strength than longer polycatenanes. For example, a 5-ring polycatenane shows increased stretch at $\mathrm{Wi}=0.1$, but is delayed to $\mathrm{Wi}=0.2$ for a 65-ring polycatenane. For $10^{0}<\mathrm{Wi}<\ 10^{2}$, there is a regime with increasing extension, which corresponds to chains being extended in flow direction with corresponding thinning in the gradient and cross-flow directions, shown in SI Figs. B.3a and B.3b, respectively. In chains with 21 or more rings, there is a decreasing regime in the $100\leq\mathrm{Wi}$ range. This corresponds to expansion in cross-flow direction, shown in SI Fig. B3.b. This behavior could not be verified in chains with 5 and 10 rings because of failure of the mechanical bonds due to bond crossing, a subject of future work. At low to moderate Wi, $\langle X_{bb}\rangle/L$ is higher in polycatenanes than linear polymers with the same molecular weight due to entangled ring topologies leading to strong repulsive interactions, keeping the chains more extended. or example, at $\mathrm{Wi}=1$, a 10-ring polycatenane has a fractional extension of 0.32 compared to a linear polymer of the same molecular weight ($M_{w}=80$) with an extension of 0.13. At high Wi, linear polymers are more extended, due to the more pronounced decreasing regime in polycatenanes. Based on 8a, increasing the ring size has the same effect on extension at low to medium Wi as number of rings. The greater flexibility of larger rings leads to more coiling behavior and less extended configurations.
@@
Polycatenanes show significantly lower shear viscosity than linear polymers of similar molecular weight, as shown in Fig. 7b. For example at $\mathrm{Wi}=10$, the viscosity of a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) is 35% of the linear polymer with 80 beads. This decreased viscosity comes from the decreased size. Compared with linear polymers of the same molecular weight, polycatenanes have smaller size Lei et al. (2021); Pakula and Jeszka (1999), also demonstrated by equilibrium $R_{g}$ data in SI Table A.4. For example, a polycatenane with $M_{w}=80$ has $R_{g}=7.36$, about 57% of linear polymer with the same molecular weight, $R_{g}=12.85$. While their viscosity is lower, polycatenanes also have weakened shear-thinning. The viscosity power-law exponent for polycatenanes is 50% less than the linear polymers at the same molecular weight, shown in SI Table E.1. Consistent with both polyrotaxanes and daisy chains, polycatenanes also do not align as well with the flow as linear polymers, showcased by their larger orientation angles at high Wi in SI Fig. B.3c. Flow alignment is better for longer polycatenanes, leading to more shear thinning for these larger molecules. From Fig. 8b, increasing the ring size in polycatenanes leads to a higher viscosity, primarily because of the increase in polymer size. The shear thinning exponent (SI Table E.1) also increases slightly for larger rings. This is due to the additional molecular degrees of freedom, leading to improved alignment with the flow, demonstrated with decreased orientation angles in SI Fig. B.4c.
@@
At very low Wi number, short, stiff polycatenanes have decreased tumbling frequency compared to linear polymers, as shown in Fig. 7c. However, because of the interlocked rings, polycatenanes are thicker in gradient direction (as seen in SI Fig. B.3a), making them more responsive to the flow strength. As a result, a slight increase in Wi causes more tumbling in short polycatenanes compared with long polycatenanes and linear polymers, where tumbling frequency remains fairly constant in low Wi regime. For example, a polycatenane with 5 rings has increased tumbling frequency at $\mathrm{Wi}\approx 0.2$, compared to linear polymers and polycatenanes with more than 10 rings where the increase occurs around $\mathrm{Wi}\approx 1$. At moderate to high Wi number, polycatenanes tumble more frequently than linear polymers. For instance, at $\mathrm{Wi}=200$, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) tumbles 3 times more frequently than the linear polymer with 80 beads. Additionally, more rings in a polycatenane decreases tumbling frequency at moderate to high Wi number. This can be explained by the resulting increase in aspect ratio of the polymer, making the polymer thickness less effective, causing the chains to behave more like linear polymers. Increasing the ring size has a similar effect on tumbling frequency as increasing the number of rings, as shown in Fig. 8c. The larger more flexible rings cause the polycatenanes to behave more like linear polymers with decreased tumbling frequency.
@@
According to Fig. 7d, polycatenanes have a $\psi_{1}$ value about two orders of magnitude lower than that of linear polymers at the same molecular weight, with the number of rings having only a minimal effect. As shown in SI Table E.2, the power-law exponent ($\beta$) for the polycatenanes is slightly higher than linear polymers of comparable molecular weight, with more rings further increasing the exponent. For example, for a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$), $\beta=-1.067$ compared with a value of -1.239 corresponding to a linear polymer with 80 beads.
@@ IV Discussion
3 Refer to caption
Figure 9: Comparison of linear and MIP polymers of similar molecular weight in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (red) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 8 beads each $M_{w}=80$.
@@ 3 Refer to caption
Polycatenanes are more extended and have lower first normal stress difference than linear chains except at very low Wi. They lack the smooth transition from coiled to stretched states typical of linear polymers. Polycatenanes show significantly lower extensional and shear viscosities, as well as weaker shear thinning. More or larger rings in polycatenanes increases viscosity, primarily due to increased size. Polycatenanes tumble more frequently, with fewer or smaller rings enhancing this effect.
@@
Overall, the interlocked topology of each MIP leads to a distinct set of rheological properties, as illustrated in Fig. B.5. Here we give an example of each polymer type, all with a similar molecular weight. We compare polymer types with similar maximum contour lengths ($L$) in SI Fig. B.5. For a given molecular weight, the polycatenane exhibits the greatest relative stretch in extensional flow (Fig. B.5a), while the polyrotaxane shows the greatest stretch in shear flow (Fig. B.5c). In shear, the daisy chain and the polycatenane have fairly consistent stretch across a wide range of Wi, varying only between 0.18-0.31 and 0.28–0.34, respectively, much less than the variability observed in linear polymers and polyrotaxanes. For a given molecular weight, the polymers also have distinct viscosity profiles. In extensional flow, the linear polymer has the greatest extensional viscosity for all but the lowest Wi range, as seen in Fig. B.5b. The mechanical bonds in all MIP types lead to lower extensional viscosity, with the daisy chain behaving similar to the linear polymer and the polyrotaxane and polycatenane showing weaker scaling of their extensional viscosity than the linear polymer and daisy chain. In shear flow (Fig. B.5d), the linear polymer has higher shear viscosity than the MIPs except the polyrotaxane, which overtakes the linear polymer at a cross-over around $\mathrm{Wi}\approx 50$. This demonstrates the reduced shear thinning of the polyrotaxane. Indeed all of the MIPs’ shear viscosity scales less strongly with Wi than the linear polymer, demonstrating that the mechanical bonds suppress shear-thinning behavior. Similar to extensional viscosity, the shear viscosity of the polycatenane is the lowest of all polymer types.
@@
In line with the varied rheological responses that we have reported, MIPs have already shown their potential in a range of applications. Slide-ring gels, a category of polyrotaxanes in which mechanically interlocked rings function as movable cross-links, could potentially be used as protective coatings in automobiles and smartphones due to their scratch-resistance and self-healing properties Noda et al. (2014). Cyclodextrin-based polyrotaxanes have shown promise in stretchable electronics and energy storage materials. Their innovative use in stretchable conductive composite and binders for anodes in lithium-ion batteries helps to overcome critical challenges in these areas Du et al. (2024). Olympic gels (materials formed entirely from interlocked rings) Raphaël et al. (1997), recently synthesized from over 16,000 DNA rings, exhibit unique nonlinear elasticity and swelling behavior with potential applications in synthetic biology and bioengineering, including artificial micro-reactors for targeted delivery and artificial nuclei in synthetic cells Speed et al. (2024). Moreover, mechanically interlocked [c2]daisy chain structures embedded in polymer networks enable advanced shape-memory materials for use as actuators in releasing and lifting objects, with further potential in soft robots via programmed stimuli-responsive motions Zhou et al. (2024).
@@
Our results hint at additional potential uses of MIPs, as well. High molecular weight polymers are known to reduce drag in pipe flow Larson (2003); Perkins et al. (1997); Dimitropoulos et al. (1998); Paterson and Abernathy (1970), at least partially due to the suppression of vortex formation caused by extended polymer configurations. The lower shear viscosity for MIPs compared to typical linear polymers and the greater extension of polyrotaxanes, in particular, suggest that MIPs could be developed into drag reducing agents. Similarly, extended polymers in liquid petroleum-based fuels favor the formation of larger droplets that are less prone to ignition during accidental fuel releases Chao et al. (1984); Wei et al. (2015); Lhota et al. (2024). The greater extension of MIPs also suggests their potential development as mist-suppressing fuel additives to improve liquid fuel safety.
@@
As a first step toward characterizing the flow behavior of mechanically interlocked polymers, this work provides valuable insights into MIP dynamics in dilute solutions under weak to moderate flow conditions. However, strong flow regimes remain unexplored as our current model fails to maintain polymer integrity at high Wi. Under such conditions, large hydrodynamic forces can cause the rings in all three MIP architectures to cross the polymer backbone, or each other in the case of polycatenanes. Although we could change the model parameters to prevent bond crossing at very high Wi, we chose to use the standard set of FENE and WCA parameters to ensure that our results remain directly comparable with previous studies in the low to moderate Wi regime. Despite the model deficiencies at high Wi, the observed degradation is nevertheless consistent with experimental findings, which show that rotaxanes degrade via dethreading and unstoppering under strong mechanical forces Muramatsu et al. (2021); Zhang and De Bo (2019). Other experimental observations suggest that mechanical bonds in catenanes are comparable in strength to covalent bonds Lee et al. (2016). To fully capture degradation mechanisms in strong flow, future work will move beyond the FENE bond model and incorporate dissociable bonds, revealing combined chemical and mechanical modes of polymer failure. This future work will lead to a better understanding of the durability of MIPs in flow for potential engineering applications.

### v3

@@ III.5 Polycatenanes: Extensional Flow
Fig. 6a shows that the variation of fractional extension of polycatenanes in the flow direction is less than linear polymers, and it is lower in polycatenanes that have fewer rings. For instance, a polycatenane with 8 beads per ring and 5 rings extends only 2.4 times beyond equilibrium, while a polycatenane chain with the same ring size and 65 rings stretches 4.2 times. More rings give the chain more degrees of freedom, leading to greater relative fluctuations in size and shape. In the limiting case, if we continue increasing the number of rings indefinitely, the chain approaches a linear polymer brereton2001statistical. At high Wi, the difference becomes smaller as both polycatenanes and linear polymers approach fully extended configurations. As shown in Fig. 6b, increasing the ring size while keeping the number of rings constant also leads to a more pronounced coil-stretch transition due to the higher conformational degrees of freedom from the larger rings. The chains with larger rings have more flexibility rauscher2020thermodynamics and can collapse and extend more readily. As a result, in weak flows, chains with smaller rings exhibit higher extension, whereas in strong flows, chains with larger rings show greater extension. No extension data is presented at $\mathrm{Wi}>40$ because strong hydrodynamic forces cause bond crossing between interlocked rings in this regime.
@@
Extensional viscosity ($\eta_{e}$) is significantly lower in polycatenanes compared to linear polymers with the same molecular weight, as shown in Fig. 6e. For instance, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) has 22 times lower ($\eta_{e}$) than a linear polymer with 80 beads at $\mathrm{Wi}=10$. Furthermore, adding more rings to a polycatenane increases the extensional viscosity due to increased molecular weight. The same effect is achieved by making the rings larger, as seen in Fig. 6f. Comparing a polycatenane with 16 beads per ring and 10 rings ($M_{w}=160$) to a polycatenane with 8 beads per ring and 21 rings ($M_{w}=168$) allows us to decouple the effects of molecular weight from the impact of chain topology on viscosity. Based on Figs. 6e and 6f, at low Wi, the chain with more small rings has higher viscosity. This reflects the more extended conformation of the smaller rings, leading to higher stress and resistance to extension. At high Wi, the chain with fewer large rings show significantly higher extensional viscosity, up to 270% higher at the highest Wi. Under strong extensional flow, the rings in the chain with fewer large rings are more effectively stretched, leading to an increased resistance to further deformation. The extended conformation under high strain produces a larger stress response, which is reflected in the higher measured extensional viscosity. In contrast, the greater number of flexible mechanical interlocks in the chain with more small rings facilitates lower stress response when subjected to high strain rates which ultimately reduces $\eta_{e}$ relative to the chain with fewer large rings. Previous polycatenane stretching simulations identified a stress-softening regime driven by rotational sliding of rings under high force chen2024nonlinear. We observed no equivalent decrease in extensional viscosity; this likely stems from differences in topology and deformation mode (constant force stretching versus extensional flow). However, we cannot rule out this behavior at high Wi number regime, where bond crossing limits our data availability.
@@ III.6 Polycatenanes: Shear Flow
As shown in Fig. 7a, at low Wi, polycatenanes exhibit a range of extensional behaviors based on the number of rings in their structure. Chains with fewer rings have limited variation of fractional extension throughout the Wi range, while chains with higher number of rings can be deformed more readily in the flow, leading to larger variation of fractional extension in flow direction. While it is somewhat trivial that smaller polymers stretch less than longer polymers, the scaling of catenanes is different than that of linear polymers. For a linear polymer in good solvent in equilibrium, the radius of gyration scales at $N^{0.6}$ de1972exponents and the contour length as $N$. This leads to a ratio of maximum stretch to equilibrium extent that scales as $N^{1.67}$. We find that from fitting equilibrium data for polycatenanes in SI Table A.4, the radius of gyration scales as $N^{0.65}$ and from our pulling simulations to determine contour length, the maximum stretch scales as $N^{0.95}$. This leads to a ratio of stretch to coil of $N^{1.46}$ for polycatenanes, demonstrating that they will stretch less than linear polymers. The extensional data in Fig. 7a reflects this resistance to stretching. For example, a polycatenane of 5 rings stretches between 0.32 and 0.43 of its maximum, compared to a 65 ring polycatenane, which extends between 0.17 and 0.35. Shorter, less flexible polycatenanes begin to stretch in the flow direction at lower flow strength than longer polycatenanes. For example, a 5-ring polycatenane shows increased stretch at $\mathrm{Wi}=0.1$, but is delayed to $\mathrm{Wi}=0.2$ for a 65-ring polycatenane. For $10^{0}<\mathrm{Wi}<\ 10^{2}$, there is a regime with increasing extension, which corresponds to chains being extended in flow direction with corresponding thinning in the gradient and vorticity directions, shown in SI Figs. B.3a and B.3b, respectively. In chains with 21 or more rings, there is a decreasing regime in the $100\leq\mathrm{Wi}$ range. This corresponds to expansion in vorticity direction, shown in SI Fig. B3.b. This behavior could not be verified in chains with 5 and 10 rings because of failure of the mechanical bonds due to bond crossing, a subject of future work. At low to moderate Wi, $\langle X\rangle/L$ is higher in polycatenanes than linear polymers with the same molecular weight due to entangled ring topologies leading to strong repulsive interactions, keeping the chains more extended. or example, at $\mathrm{Wi}=1$, a 10-ring polycatenane has a fractional extension of 0.32 compared to a linear polymer of the same molecular weight ($M_{w}=80$) with an extension of 0.13. At high Wi, linear polymers are more extended, due to the more pronounced decreasing regime in polycatenanes. Based on 8a, increasing the ring size has the same effect on extension at low to medium Wi as number of rings. The greater flexibility of larger rings leads to more coiling behavior and less extended configurations.
@@
Polycatenanes show significantly lower shear viscosity than linear polymers of similar molecular weight, as shown in Fig. 7b. For example at $\mathrm{Wi}=10$, the viscosity of a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) is 35% of the linear polymer with 80 beads. This decreased viscosity comes from the decreased size. Compared with linear polymers of the same molecular weight, polycatenanes have smaller size lei2021dimensional; pakula1999simulation, also demonstrated by equilibrium $R_{g}$ data in SI Table A.4. For example, a polycatenane with $M_{w}=80$ has $R_{g}=7.36$, about 57% of linear polymer with the same molecular weight, $R_{g}=12.85$. While their viscosity is lower, polycatenanes also have weakened shear-thinning. The viscosity power-law scaling exponent can be up to 50% less than the linear polymers, shown in SI Table E.1. For example, compared to the scaling exponent of -0.38 for a linear polymer of $M_{w}=80$, the 10-ring, 8-bead-per-ring polycatenane has a much reduced scaling of -0.154. Generally, the scaling of polycatenanes is reduced from linear polymers with shorter polycatenanes or smaller rings leading to weakened shear thinning. Consistent with both polyrotaxanes and daisy chains, polycatenanes also do not align as well with the flow as linear polymers, showcased by their larger orientation angles at high Wi in SI Fig. B.3c. Flow alignment is better for longer polycatenanes, leading to more shear thinning for these larger molecules. From Fig. 8b, increasing the ring size in polycatenanes leads to a higher viscosity, primarily because of the increase in polymer size. The shear thinning exponent (SI Table E.1) also increases slightly for larger rings. This is due to the additional molecular degrees of freedom, leading to improved alignment with the flow, demonstrated with decreased orientation angles in SI Fig. B.4c.
@@
At very low Wi number, short, stiff polycatenanes have decreased tumbling frequency compared to linear polymers, as shown in Fig. 7c. However, because of the interlocked rings, polycatenanes are thicker in gradient direction (as seen in SI Fig. B.3a), making them more responsive to the flow strength. As a result, a slight increase in Wi causes more tumbling in short polycatenanes compared with long polycatenanes and linear polymers, where tumbling frequency remains fairly constant in low Wi regime. For example, a polycatenane with 5 rings has increased tumbling frequency at $\mathrm{Wi}\approx 0.2$, compared to linear polymers and polycatenanes with more than 10 rings where the increase occurs around $\mathrm{Wi}\approx 1$. At moderate to high Wi number, polycatenanes tumble more frequently than linear polymers. For instance, at $\mathrm{Wi}=200$, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) tumbles 3 times more frequently than the linear polymer with 80 beads. Additionally, more rings in a polycatenane decreases tumbling frequency at moderate to high Wi number. This can be explained by the resulting increase in aspect ratio of the polymer, making the polymer thickness less effective, causing the chains to behave more like linear polymers. Increasing the ring size has a similar effect on tumbling frequency as increasing the number of rings, as shown in Fig. 8c. The larger more flexible rings cause the polycatenanes to behave more like linear polymers with decreased tumbling frequency. SI Table E.3 shows that polycatenanes with 8-bead rings have a much stronger scaling of $\omega\tau_{R}$ with Wi than linear polymers. This scaling increases with more rings. For example, a 5-ring polycatenane with 8-bead rings exhibits a scaling exponent of 0.917 with tumbling frequency and a 65-ring polycatenane with 8-bead rings has an exponent of 1.029. In both cases the scaling is significantly increased from the expected linear polymer scaling of 0.75. Conversely, increasing the ring size reduces the exponent toward the linear polymer value. For example, while a 10-ring polycatenane with 8-bead rings has a scaling exponent of 0.935, a 32-ring polycatenane with the same size rings has an exponent of 0.753, identical to that expected from a linear polymer.
@@
According to Fig. 7d, polycatenanes have a $\psi_{1}$ value about two orders of magnitude lower than that of linear polymers at the same molecular weight, with the number of rings having only a minimal effect. As shown in SI Table E.2, the power-law scaling exponent ($\beta$) for the polycatenanes is slightly higher than linear polymers of comparable molecular weight, with more rings further increasing the exponent. For example, for a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$), $\beta=-1.067$ compared with a value of -1.239 corresponding to a linear polymer with 80 beads.
@@ IV Discussion
Figure 9: Comparison of linear and MIP polymers of similar molecular weight in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction $(\langle X\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (red) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 8 beads each ($M_{w}=80$).
@@
Polycatenanes are more extended and have lower first normal stress difference than linear chains except at very low Wi. In line with previous experimental work on kinetoplasts (catenated DNA networks) soh2020deformation, these polymers do not show an abrupt transition from coiled to stretched states typical of linear polymers. Polycatenanes show significantly lower extensional and shear viscosities, as well as weaker shear thinning. More or larger rings in polycatenanes increases viscosity, primarily due to increased size. Polycatenanes tumble more frequently, with fewer or smaller rings enhancing this effect.
@@
Overall, the interlocked topology of each MIP leads to a distinct set of rheological properties, as illustrated in Fig. B.5. Here we give an example of each polymer type, all with a similar molecular weight. For direct comparison with a linear polymer, properties of MIPs are normalized by those of a linear polymer, as shown in SI Fig. B.6. We compare polymer types with similar maximum contour lengths ($L$) in SI Fig. B.5. For a given molecular weight, the polycatenane exhibits the greatest relative stretch in extensional flow (Fig. B.5a), while the polyrotaxane shows the greatest stretch in shear flow (Fig. B.5c). In shear, the daisy chain and the polycatenane have fairly consistent stretch across a wide range of Wi, varying only between 0.18-0.31 and 0.28–0.34, respectively, much less than the variability observed in linear polymers and polyrotaxanes, showing a weaker coil–stretch response. Additionally, similar to the effect of knots on relaxation dynamics and coil-stretch transition of DNA molecules soh2018knots, mechanical bonds decrease the relaxation time of the MIPs (see SI Tables A.1 to A.4) by lowering the contour length of the polymer, and shift the coil-stretch transition to higher strain rates. For a given molecular weight, the polymers also have distinct viscosity profiles. In extensional flow, the linear polymer has the greatest extensional viscosity, as seen in Fig. B.5b. The mechanical bonds in all MIP types lead to lower extensional viscosity, with the daisy chain behaving similar to the linear polymer and the polyrotaxane and polycatenane showing weaker scaling of their extensional viscosity than the linear polymer and daisy chain. In shear flow (Fig. B.5d), the linear polymer has higher shear viscosity than the MIPs except the polyrotaxane, which overtakes the linear polymer at a cross-over around $\mathrm{Wi}\approx 50$. This demonstrates the reduced shear thinning of the polyrotaxane. Indeed all of the MIPs’ shear viscosity scales less strongly with Wi than the linear polymer, demonstrating that the mechanical bonds suppress shear-thinning behavior. This suppression is further reflected in the increase of the normalized viscosities shown in SI Fig. B.6d. Similar to extensional viscosity, the shear viscosity of the polycatenane is the lowest of all polymer types.
@@
In line with the varied rheological responses that we have reported, MIPs have already shown their potential in a range of applications. Slide-ring gels, a category of polyrotaxanes in which mechanically interlocked rings function as movable cross-links, could potentially be used as protective coatings in automobiles and smartphones due to their scratch-resistance and self-healing properties noda2014topological. Cyclodextrin-based polyrotaxanes have shown promise in stretchable electronics and energy storage materials. Their innovative use in stretchable conductive composite and binders for anodes in lithium-ion batteries helps to overcome critical challenges in these areas du2024cyclodextrins. Olympic gels (materials formed entirely from interlocked rings) raphael1997progressive, recently synthesized from over 16,000 DNA rings, exhibit unique nonlinear elasticity and swelling behavior with potential applications in synthetic biology and bioengineering, including artificial micro-reactors for targeted delivery and artificial nuclei in synthetic cells speed2024assembling. Moreover, mechanically interlocked [c2]daisy chain structures embedded in polymer networks enable advanced shape-memory materials for use as actuators in releasing and lifting objects, with further potential in soft robots via programmed stimuli-responsive motions zhou2024mechanically.
@@
Our results hint at additional potential uses of MIPs, as well. High molecular weight polymers are known to reduce drag in pipe flow larson2003analysis; perkins1997single; dimitropoulos1998direct; paterson1970turbulent, at least partially due to the suppression of vortex formation caused by extended polymer configurations. The lower shear viscosity for MIPs compared to typical linear polymers and the greater extension of polyrotaxanes, in particular, suggest that MIPs could be developed into drag reducing agents. Similarly, extended polymers in liquid petroleum-based fuels favor the formation of larger droplets that are less prone to ignition during accidental fuel releases chao1984antimisting; wei2015megasupramolecules; lhota2024mist. The greater extension of MIPs also suggests their potential development as mist-suppressing fuel additives to improve liquid fuel safety.
As a first step toward characterizing the flow behavior of mechanically interlocked polymers, this work provides valuable insights into MIP dynamics in dilute solutions under weak to moderate flow conditions. However, strong flow regimes remain unexplored as our current model fails to maintain polymer integrity at high Wi. Under such conditions, large hydrodynamic forces can cause the rings in all three MIP architectures to cross the polymer backbone (polyrotaxanes), or each other (daisy chains and polycatenanes). In extensional flow, these bond crossings occur at flow strengths less than those required to fully elongate the MIPs. As a result, the extension curves do not saturate to 1 for this model (Figs. 2a, 4a, 6a, b). Moreover, shielding from hydrodynamic interactions (HI) substantially increases the Wi at which bond crossing occurs. To isolate this effect, we simulated the polymers shown in Fig. 9 without hydrodynamic interactions; the results are provided in SI Fig. B.7. In the absence of HI, bond crossing happens at significantly lower Wi, as expected when the shielding effect of HI is removed and interior beads are subject to full hydrodynamic forces. Correspondingly, all polymer types stretch more in both extensional and shear flow and exhibit higher viscosities. The shear thinning trends and scaling, however, are preserved in the absence of HI.
@@
Although we could change the model parameters to prevent bond crossing at very high Wi, we chose to use the standard set of FENE and WCA parameters to ensure that our results remain directly comparable with previous studies in the low to moderate Wi regime. Despite the model deficiencies at high Wi, the observed degradation is nevertheless consistent with experimental findings, which show that rotaxanes degrade via dethreading and unstoppering under strong mechanical forces muramatsu2021rotaxane; zhang2019mechanical. Other experimental observations suggest that mechanical bonds in catenanes are comparable in strength to covalent bonds lee2016mechanical. To fully capture degradation mechanisms in strong flow, future work will move beyond the FENE bond model and incorporate dissociable bonds, revealing combined chemical and mechanical modes of polymer failure. This future work will lead to a better understanding of the durability of MIPs in flow for potential engineering applications.